1. Suppose that, for the production process for quarks, labour and capital are gross complements. Using isoquant / isocost analysis, illustrate how the long-run quantity-demanded for labour would be affected by a decrease in the cost of capital. Illustrate both the “substitution effect” and the “scale effect.”

2. Suppose capital and labour are perfect substitutes for producing widgets and that it takes 1 unit of capital or 3 units of labour to produce one widget. If the price of labour is $4 and the price of capital is $16, how much labour and capital should a firm use to produce 30 widgets? Explain.

3. All else equal, will an increase in the productivity of capital always decrease the quantity of labour that a firm hires? Illustrate and explain using isoquant / isocost analysis.

4. With appropriate illustrations and explanations, use the concept of marginal revenue product to explain how, all else equal, the demand for labour of a competitive market differs from a monopoly.

5. Which is more elastic, the demand for pharmacists or the demand for custodial workers. Discuss with reference to any two of the Laws of Derived Demand.
Answer #1

Due to presence of HOMEWORKLIB POLICY, I am answering one question.

2.

Ans:

Output produced using 1 extra unit of capital = 1 = MPK

Output produced using 1 extra unit of labor = 1/3 = MPL

MRTS = MPL / MPK = 1 / 3

Price ratio = 4 / 16 = 1/4

Since MRTS > Price ratio, marginal product per cost is greater in labor.

Hence, only labor should be used.

Labor = 30 X 3 = 90 units.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!